Find the closure of the following set in C

{z∈C : (ReZ)2=ImZ}


1
Expert's answer
2021-08-19T11:29:31-0400

The set is a curve in "\mathbb{C}". Thus, it is closed, since the set coincides with its boundary. Since the set is closed, the closure of the set is the set itself, i.e., "\{z\in{\mathbb{C}}:(Re z)^2=Im z\}".